OpenStudy (anonymous):

can someone show me work? Can someone please help? Solve the system by substitution. -4.5x-2y=-12.5 3.25x-y=-0.75

OpenStudy (anonymous):

Multiply the second equation by 2: (3.25x-y=-0.75) x 2 =6.5x -2y=-1.5 Now subtract this new equation from the first one to eliminate y, i.e: -4.5x-2y=-12.5 -(6.5x -2y=-1.5) =>-11x+0=-11 x=-11/-11=1 put x back into the first eqn: -4.5(1)-2y=-12.5 -4.5-2y=12.5 -2y=12.5+4.5 y=-8.5 so, x=1, y=8.5.

OpenStudy (anonymous):

thank you so much!! I relly appreciate it!! (:

OpenStudy (anonymous):

you're welcome! :)

OpenStudy (anonymous):

wait when you subtract the -2 from the two wouldnt you get 4?

OpenStudy (anonymous):

the -2y - 2y isnt that 4?

OpenStudy (anonymous):

No, it's -2y-(-2y)=-2y+2y=0

OpenStudy (anonymous):

ohhhh i see what you mean! Im sorry i got confused for a second! thank you again!

OpenStudy (anonymous):

that's alright :)